(2) Solow Model Arithmetic: Suppose that the economy has the following production function: K > 0 n > The population grows at the exogenously given rate n, so that N,-(1 + n) (a) Derive the per worker production function, where y - Y/N is output per worker and k- K/N is capital per worker (b) Derive the aggregate accumulation equation for capital per worker expressed solely as a function of k, k, A. and parameters (s, θ, d, n). Recall the law of motion for capital: C=(1-s)y (c) Show that this accumulation equation is strictly concave (i.e. it has a second derivative with respect to k that is strictly less than 0 (d) Define the steady state (in words) e) Derive an expression for the steady state capital stock per work as a function of A, d, s, θ, and n. (f) Derive an expression for the steady state consumption per worker. (g) Find the saving rate, s, that maximizes steady state consumption per worker. To do this, take the derivative of your answer to part (d) with respect to s, set it equal to zero and simplify, finding an expression for s as a function of the other parameters of the
Answer #1

a ] the production function is given as

y = A K \Theta N1-\Theta

The per worker term is given as

y = y/N = A K\Theta N1-\Theta/ N = A ( K/N)\Theta = AK\Theta

b) the aggregate accumulation equation is

k / = (1-d) K +i

=(1-d) K+ s y

per worker term ; k //N / = (1-d) k+ s y / N /

= (1-d) k + s y / (1+n) N

K / = 1-d / 1+n k + s/1+n y

= 1-d/ 1+n k + s / 1+n A K \Theta

C)  the second derivatives of k / is ; d 2 k / / d k 2 =d/ dk [ dk/ / dk ]

= d/dk [ 1-d / 1+n + SA\Theta / 1+n k \Theta - 1 ]

= SA\Theta(\Theta -1 ) / 1+n k \Theta - 2

AS  \Theta < 1 ; \Theta - 1 < 0 ;

d /  SA\Theta(\Theta-1) / 1+N K\Theta - 2 < 0

d 2  k / / d k 2 < 0 [ proved ]

d) At the study the capital per worker is constant and output grows at a constant rate

e) At the steady state ; k / = k = k *

= k * = 1-d / 1+n k* + S A K* \Theta /1+n

(n+d*) k * S A K* \Theta

K* = (SA/ n+d) 1/ 1-\Theta

f) the consumption per worker is given as

c * = (1-s) y *

= (1-s) [ SA / n+d ] \Theta/ 1-\Theta = (SA/n+d) \Theta/1-\Theta - S 1/1-\Theta(A/n+d) \Theta/1-\Theta

g) at the maximum c * ;

\partialc* / \partials = (A / n+d ) \Theta/1-\Theta[ \Theta/ 1- \Theta s 2\Theta-1/1-\Theta- 1/ 1-\Theta s \Theta/ 1-\Theta ] = 0

\Theta/ 1-\Theta s 2\Theta-1 / 1-\Theta  = 1/ 1- \Theta s \Theta/ 1-\Theta

or s2\Theta-1 / 1-\Theta -\Theta/ 1-\Theta =1/ \Theta

or s\Theta-1/1-\Theta= 1/ \Theta

or 1/s = 1/\Theta

\therefore s * = \Theta